OpenStudy (anonymous):

Solve the proportion 20/100 = x/147.6 Find the answer to the nearest tenth. Please explain your answer.

OpenStudy (anonymous):

\[\frac{ 20 }{ 100 }=\frac{ x }{ 147.6 }\]

OpenStudy (anonymous):

Cross multiply.

OpenStudy (anonymous):

I see now! thank you!! duh, I feel stupid now :P

OpenStudy (anonymous):

OK, glad I could sort of help you.

OpenStudy (anonymous):

you were a big help! thanks!

OpenStudy (anonymous):

\[100x=2952\] \[100x/100=x\] \[2952/100=29.52\] \[x=29.52\] I just worked it out anyways. It is nice to get practice every once and a while to sharpen your brain.